OpenStudy (anonymous):

To convert from degrees to radians, multiply by 180degrees/pi? True or False

OpenStudy (solomonzelman):

it is actually times pi/180degrees though....

OpenStudy (solomonzelman):

So I guess it is false.

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks u buddy

OpenStudy (solomonzelman):

Yes, but the question says T/F whether it is times 180/pi, and it is false.

OpenStudy (solomonzelman):

Becuase the truth is, times pi/180 degrees

OpenStudy (anonymous):

yeah its times so its false!

OpenStudy (anonymous):

notice\[\frac{ 180^{\circ} }{ \pi \text{ radians} }\ne\frac{ \pi \text{ radians} }{ 180^{\circ} }\]

OpenStudy (anonymous):

same conversion factor but used to get different things. if you want to convert degrees to radians, use the one on the right. if you want to convert radians to degrees, use the one on the left